OpenStudy (anonymous):

Write the slope-intercept form of the equation of the line through the given points. Show your work. Through: (2,-2) (-5,5)

OpenStudy (anonymous):

Do you know what the slope-intercept form is?

OpenStudy (anonymous):

Yeah but I got to -2=-3/7920+b I don't know what to do next @faith2012

OpenStudy (anonymous):

I meant -2=-3/7 (2) +b

OpenStudy (anonymous):

@Shadowgirl15

OpenStudy (anonymous):

You solve it now.

OpenStudy (anonymous):

But how do I multiply 2 by the -3/7? @faith2012

OpenStudy (anonymous):

-3/7*2/1

OpenStudy (anonymous):

Multiply straight across.

OpenStudy (anonymous):

\[\frac{ -3 }{ 7 } * \frac{ 2 }{ 1 }\] =?

OpenStudy (anonymous):

That'll give me -6/7 then I subtract that from -2 right? Do I make -2 a fraction to?

OpenStudy (anonymous):

@faith2012

OpenStudy (anonymous):

You'd add 2 to both sides, and yes. Too make it easier, you'd make it into a fraction with the same denominator as -6/7

OpenStudy (anonymous):

Then that means it'll look like this? 2+-6/7 -> 2/7+-6/7=14/7+-6/7=8/7 (?) @faith2012

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Good job! :)

OpenStudy (anonymous):

Thank you! @faith2012

OpenStudy (anonymous):

I'll give you a medal because you did really great!

OpenStudy (anonymous):

Aww~ Thanks :D @faith2012

OpenStudy (anonymous):

You're welcome :)